1) Calculate the molarity of a 45.3 g NaNO3 dissolved in water to produce 225 mL of solution.



2) What mass of concentrated HCl is needed to prepare 2.45 L of 13.0 M HCl?



3) What volume can be prepared from 1.0 M NaOH from 2.54 g NaOH?

1
Expert's answer
2022-03-17T14:58:44-0400

1)Molarity= moles/ litre


Moles= 45.3/85= 0.5329

Molarity= 0.53M

2)mass= moles× rmm

Moles=( 2.45/1)×13= 31.85

Mass= 31.85× 36.5= 1162.5g

3)moles= 2.54/ 40= 0.0635

If 1.0 moles are in a litre then

(0.0635/1)× 1000= 63.5ml